During a baseball game, a batter hits a high
pop-up.
If the ball remains in the air for 5.26 s, how
high does it rise? The acceleration of gravity
is 9.8 m/s
1
Expert's answer
2019-09-27T09:34:17-0400

The time of motion of the ball up while it reaches a maximum height


"t=\frac{5.26\:\rm s}{2}=2.63\: \rm s"

The maximum height


"h=\frac{gt^2}{2}=\frac{9.8\:\rm m/s^2\times (2.63\:\rm s)^2}{2}=33.9\:\rm m"